OpenStudy (anonymous):

Simplify the following expression. 3x − 6y + 2z − 5x − 3y + 8z −2x + 9y + 10z −2x + 9y − 10z −2x − 9y + 10z 2x − 9y + 10z

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):



OpenStudy (anonymous):

yup

OpenStudy (anonymous):

then what

OpenStudy (anonymous):

once you combin like terms you will get one of ur options

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

3x+(-5x)=-2x -6y+(-3y)=-9y 2z+8z=10z

OpenStudy (anonymous):

-2x+-9y+10z

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

Evaluate: 4 − 35 ÷ 7 + 32 −10 4.6 8 5

OpenStudy (anonymous):

* 3^2

OpenStudy (anonymous):

No I would be -2x-9y+10z

OpenStudy (anonymous):

yea same thing how about the other one Evaluate: 4 − 35 ÷ 7 + 3^2 −10 4.6 8 5

OpenStudy (anonymous):

4-35/7+9 4-5+9 4-14 10

OpenStudy (anonymous):

i got 8

OpenStudy (anonymous):

could you show me what you did to get that

OpenStudy (nirmalnema):

its the 3 one correct...add all the same terms..

OpenStudy (anonymous):